OpenStudy (anonymous):

Parabolas! axis of symmetry at x=-7 a maximum height of 4 and also passes through the point (-6,0) write the equation of the parabola in vertex form, will award a medal and fan !

OpenStudy (jjuden):

The vertex form of the equation is y = a(x - h)2 + k

OpenStudy (anonymous):

okay so what would be a, h , and k?

OpenStudy (anonymous):

You have to solve for a by plugging in things you know. (h,k) is the vertex, (x,y) can be any point you have. axis of symmetry is also the x coordinate of the vertex, maximum height is the y value of the vertex so you have a vertex of (-7,4) so plugging in things we know... vertex (-7,4) point (-6,0) 0=a(-6-(-7))+4 can you solve for a?

OpenStudy (anonymous):

Forgot my squared... 0=a(-6-(-7))^2+4

OpenStudy (jjuden):

Since the axis of symmetry is a vertical line, x = -7 and the parabola has a maximum height of 4 units above the x axis; we can safely conclude that it opens down and its vertex is at (-7, 4). Besides, it passes through the point (-6, 0), which is below the vertex; definitively, it opens down. Furthermore, if it would open up, it would have a minimum height, not a maximum; so, no arguments, it opens down. The vertex form of the equation is y = a(x - h)2 + k Since the vertex is defined as (h, k), h = -7 and k = 4: y = a(x - (-7))2 + 4 or: y = a(x + 7)2 + 4 We just need to find the value of "a" (which must be negative, remember, it opens down). Here is where we use the point through which the parabola passes through: (-6, 0). We substitute the values x = -6, y = 0 in the equation and solve for a: 0 = a(-6 + 7)2 + 4 simplify and subtract 4 on both sides of the equation: -4 = a(1)2 or a = -4 Now we are ready to write our completed equation: y = -4(x + 7)2 + 4

OpenStudy (anonymous):

okay wow thank you both can I like give one of you a medal and the other gives the other one?

OpenStudy (jjuden):

sure